I worked 800o hors . and was paid 5000 dollers how much did i make per hour
Answer by fcabanski(1391)

5000/800 = 50/8 = $6.25 per hour.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)